paul is 4 years younger than his brother. in 4 years, paul will be three-fourths as old as his brother. how old are they now?
Answer by Maths68(1474)

Let
Present age of Paul = p
Present age of his brother = b


Paul is 4 years younger than his brother
p=b-4...........(1)


In 4 years,
Paul will be = p+4
His brother be = b+4
Then
Paul will be three-fourth as old as his brother
p+4=3/4(b+4)
4(p+4)=3(b+4)
4p+16=3b+12
4p-3b=12-16
4p-3b=-4
Put the value of p from (1) to above equation
4(b-4)-3b=-4
4b-16-3b=-4
4b-3b=-4+16
b=12
Put the value of b in (1)
p=b-4
p=12-4
p=8


Present age of Paul = p = 8
Present age of his brother = b = 12